OpenStudy (anonymous):

PLZ HELP!!!!! A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? Choose all answers that are correct.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

A. -3X+Y=-5 B. 3X+Y=-5 C. (Y-8)=3(X-1) D. (Y+8)=3(X+1)

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

i cant say the answer but ill say this y=mx + b where m is the slope and y is the b

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\]

pooja195 (pooja195):

m=slope (x1,y1)

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

so C is one of them ?

pooja195 (pooja195):

@pamela.g direct answers are not allowed here .

pooja195 (pooja195):

\[\huge~y-8=3(x+1)\]

pooja195 (pooja195):

i'll leave you with that.

OpenStudy (anonymous):

how is that not a direct answer? wtf @pooja195

pooja195 (pooja195):

I plugged in the info for them i did NOT give the answer. And refrain from using profanity.

OpenStudy (anonymous):

and your answer's wrong btw. oh also i'll do whatever I please @pooja195

pooja195 (pooja195):

Whatever you desire ^_^

pooja195 (pooja195):

@leutz13 y−8=3(x+1) solve that and you ahve your answer

OpenStudy (anonymous):

wt fudge how do i solve that??????

OpenStudy (anonymous):

isn't it (y+8) = 3 (x+1) ? @pooja195

OpenStudy (misty1212):

there is nothing to solve

OpenStudy (misty1212):

answer D has the form you get from a direct substitution \[y+8=3(x+1)\] there is nothing left to "solve" for

pooja195 (pooja195):

oops i wasnt looking at the answer choices lol sorry about that.....but still on the right track

OpenStudy (misty1212):

@leutz13 you can solve it for \(y\) which is what it being said here, but if you look at your answer choices you do not have to do that

OpenStudy (mathmate):

Use what @pooja195 gave you: \(\Large y−y1=m(x−x1)\) y1=-8 x1=-1 m=3 so \(\Large y-(-8) = 3(x-(-1)\) continue to get your answer.

OpenStudy (anonymous):

yeah i no one of the answers i just cant figure out the other